In 2020, Lisa and Fred, a married couple, had a taxable income of $306,700. If they were to file separate tax returns, Lisa would have reported taxable income of $127,200 and Fred would have reported taxable income of $179,500.



What is the couple’s marriage penalty or benefit? (Do not round intermediate calculations.)

Use the Tax Rate Schedule for reference. 2020 Tax Rate Schedules

Individuals

Schedule X-Single

If taxable income is over: But not over: The tax is:
$          0 $    9,875 10% of taxable income
$    9,875 $ 40,125 $987.50 plus 12% of the excess over $9,875
$ 40,125 $ 85,525 $4,617.50 plus 22% of the excess over $40,125
$ 85,525 $163,300 $14,605.50 plus 24% of the excess over $85,525
$163,300 $207,350 $33,271.50 plus 32% of the excess over $163,300
$207,350 $518,400 $47,367.50 plus 35% of the excess over $207,350
$518,400 $156,235 plus 37% of the excess over $518,400

Schedule Y-1-Married Filing Jointly or Qualifying Widow(er)

If taxable income is over: But not over: The tax is:
$          0 $ 19,750 10% of taxable income
$ 19,750 $ 80,250 $1,975 plus 12% of the excess over $19,750
$ 80,250 $171,050 $9,235 plus 22% of the excess over $80,250
$171,050 $326,600 $29,211 plus 24% of the excess over $171,050
$326,600 $414,700 $66,543 plus 32% of the excess over $326,600
$414,700 $622,050 $94,735 plus 35% of the excess over $414,700
$622,050 $167,307.50 plus 37% of the excess over $622,050

Schedule Z-Head of Household

If taxable income is over: But not over: The tax is:
$          0 $ 14,100 10% of taxable income
$ 14,100 $ 53,700 $1,410 plus 12% of the excess over $14,100
$ 53,700 $ 85,500 $6,162 plus 22% of the excess over $53,700
$ 85,500 $163,300 $13,158 plus 24% of the excess over $85,500
$163,300 $207,350 $31,830 plus 32% of the excess over $163,300
$207,350 $518,400 $45,926 plus 35% of the excess over $207,350
$518,400 $154,793.50 plus 37% of the excess over $518,400

Schedule Y-2-Married Filing Separately

If taxable income is over: But not over: The tax is:
$          0 $    9,875 10% of taxable income
$    9,875 $ 40,125 $987.50 plus 12% of the excess over $9,875
$ 40,125 $ 85,525 $4,617.50 plus 22% of the excess over $40,125
$ 85,525 $163,300 $14,605.50 plus 24% of the excess over $85,525
$163,300 $207,350 $33,271.50 plus 32% of the excess over $163,300
$207,350 $311,025 $47,367.50 plus 35% of the excess over $207,350
$311,025 $83,653.75 plus 37% of the excess over $311,025
Answer #1
Tax, if filed single:
Lisa 24607.5 =14605.5+(127200-85525)*24%
Fred 38455.5 =33271.5+(179500-163300)*32%
Total Tax, if filed single 63063
Tax, if filed jointly 61767 =29211+(306700-171050)*24%
Marriage benefit 1296 =63063-61767